Solve the polynomial inequality. Express the solution set in interval notation.
36 x^2 + 1 greaterthan or equalto 12 x

Answer by stanbon(75887)

Solve the polynomial inequality. Express the solution set in interval notation.
36 x^2 + 1 greaterthan or equalto 12x
-----
36x^2+1 >= 12x
-----
36x^2 - 12x + 1 >= 0
------
(6x-1)^2 >= 0
----
x >= 1/6 with multiplicity 2
-----
Ans: {x | x >= 1/6}
-----------
Cheers,
Stan H.
-------------------